sir, pl. help me to solve following problems of permutation & combination- Q1. how many divisors of 9000 are even but not divisible by 4?also find the sum of such nos. Q2. the no. of ways an examiner can assign 30 marks to 8 ques., awarding not less than 2 marks to any ques. is

sir,
pl. help me to solve following problems of permutation & combination-


Q1. how many divisors of 9000 are even but not divisible by 4?also find the sum of such nos.



Q2. the no. of ways an examiner can assign 30 marks to 8 ques., awarding not less than 2 marks to any ques. is

 

Grade:

1 Answers

Nishant Vora IIT Patna
9 years ago
[Q-1]9000=325323
we have to take exactly one 2.
so number of ways (2+1)(3+1)(1)= 12

Sum of such numbers =(2)(30 + 31 +32)(50 +51 +52 +53)

[Q-2] first of all give 2 marks to each student(Total 16 marks). so now we have to distribute remaining 14 marks to 8 ques and all the question can get zero or more marks. So it will be same as distributing 14 coins among 8 beggars.
So, Number of ways = 14+8-1C8-1 = 21C7

For in depth knowledge of this chapter you can refer
http://www.askiitians.com/online-study-packages.aspx

Thanks,
Nishant